OpenStudy (samantha2234):

~PLEASE HELP~ At a bargain store, Tanya bought 3 items that each cost the same amount. Tony bought 4 items that each cost the same amount, bu each was $2.25 less than the items that Tanya bought. Both Tanya and Tony paid the same amount of money. What was the individual cost of each person's items? Write an equation, and let x represent the individual cost of one of Tanya's items.

OpenStudy (3mar):

@Samantha2234 What you got? Do you know how to start?

OpenStudy (davejavous):

One of Tony's items were $2.25 less than one of Tanya's. If x is the cost of one of Tanyas items, then on of Tony's costs: \[\$ (x-2.25)\] Now 4 of Tony's items si the same price as 3 of Tanya's: \[4(x-2.25) = 3x\] Expand the bracket on the left: \[4x - 9 = 3x\] Rearrange to find x: \[4x - 3x = 9\]\[x=9\] This means that one of Tanya's items was $9.00 Tony's items are $(x-2.25), therefore: $(9-2.25) = $6.75 Tanya's Item: $9.00 Tony's Item: $6.75

OpenStudy (samantha2234):

Oh my goodness, you are a LIFE SAVER!!!!!!!!!!

OpenStudy (3mar):

Thank you, davejavous. You did it!

OpenStudy (davejavous):

Np c: If you'd give me a medal I'd really apreciate it thank you. Hope it helped.